# Regulatable ML

### v1

@@ A.2 POLICY-AS-PROMPT OUTPUT
Rules for VALID Inputs It is acceptable to automate the analysis of HR support tickets only if the content is free of personal data. Example of a VALID anonymized HR support ticket: "Request: I am unable to access the HR portal to update my benefits information. Error: System shows ’Access Denied’ despite using correct login. Resolution attempted: Cleared cache, tried on a different browser. Request: Please escalate to IT for resolution." Rules for INVALID Inputs Including any personal data in AI analysis or outputs without prior automated deletion or anonymization is a direct violation of the policy. Examples of personal data that must be removed include: names, email addresses, employee IDs, government IDs, birth dates, home addresses, salary details, beneficiary information, etc. Example of an INVALID non-anonymized HR support ticket: "Request: John Smith (Employee ID: 45832) cannot access his HR portal. Email: john.smith@company.com Issue: Unable to update dependent details for his son born on 12/03/2020. Request: Please resolve urgently." Example of an INVALID non-HR ticket (out of scope): "Request: The office coffee machine on the 3rd floor is broken again. Issue: Display shows ’Water Pump Error.’ Request: Please arrange maintenance." Output Format: Return a single JSON object with the following keys: "classification": Either "VALID" or "INVALID". "reasoning": A one-sentence explanation referencing the rule followed or violated. Task: Analyze the following DOCUMENT and provide your classification: Your JSON response:
@@
LLM Prompt: Policy Compliance Classification (Content-Focused) You are a security/policy analyst. Analyze the user’s query and decide if the content of the INPUT is allowed under the policy. Ignore tooling/functions; focus only on what data the INPUT contains and what the OUTPUT must contain. Policy — Allowed vs. Disallowed Content ALLOWED INPUT CONTENT (no personal data, no secrets) Alert metadata: alert/case IDs, timestamps, severity, source system, rule/signature names, tags. Technical indicators (IoCs): file hashes, public IPs, domains, URLs (no credentials), file/process names, registry keys, event codes. System/asset context: asset IDs, generic hostnames (not derived from real names), service/application names, environment (prod/stage/dev), non-personal ticket/channel IDs. Log details (sanitized): network flow summaries, HTTP metadata (without tokens/cookies), counts, time windows, error codes; fields may be pseudonymized or hashed. Org context (non-personal): role or team names (e.g., “SRE on-call”), runbook references, anonymized user placeholders (user_A, hashed IDs). Threat intel: malware family names, campaign/actor names, rule IDs, confidence scores. Attachments/screenshots only if redacted/pseudonymized so no personal data or secrets are visible. DISALLOWED INPUT CONTENT Personal data (PII): names, emails, phone numbers, home addresses, government IDs, birth dates, employee/customer IDs tied to individuals. Sensitive records: health, financial, or HR records; salary/beneficiary details; customer lists; resumes. Secrets/credentials: passwords, API keys, private keys, tokens, session IDs, cookies, JWTs, MFA codes, OAuth headers. Raw dumps containing PII/secrets: browser/localStorage, memory dumps, unredacted chat/email transcripts, authorization headers. Identifiers revealing a person: hostnames/usernames or filenames that embed real names (must be pseudonymized before submission). Out-of-scope facility issues (e.g., coffee machine maintenance). Output Contract Return a single JSON object: "classification": "VALID" if the INPUT only contains Allowed content; otherwise "INVALID". "reasoning": One concise sentence citing the specific rule followed/violated. Task: Analyze the following DOCUMENT and classify it: Your JSON response:
@@ A.3 Evaluation for POLICY-TREE-GEN
SecOps O1 26.9 22.6 85.7 0.987 100 0.958
SecOps GPT-OSS 120B 66.7 10.3 0.0 0.974 0.0 0.881
SecOps Llama 405B 40.0 9.8 0.0 0.842 100.0 0.790
SecOps Claude 3.5 100.0 5.4 0.0 0.818 100.0 0.782

### v2

@@ A.2 POLICY-AS-PROMPT OUTPUT
Rules for ID Inputs It is acceptable to automate the analysis of HR support tickets only if the content is free of personal data. Example of a ID anonymized HR support ticket: "Request: I am unable to access the HR portal to update my benefits information. Error: System shows ’Access Denied’ despite using correct login. Resolution attempted: Cleared cache, tried on a different browser. Request: Please escalate to IT for resolution." Rules for OOD Inputs Including any personal data in AI analysis or outputs without prior automated deletion or anonymization is a direct violation of the policy. Examples of personal data that must be removed include: names, email addresses, employee IDs, government IDs, birth dates, home addresses, salary details, beneficiary information, etc. Example of an OOD non-anonymized HR support ticket: "Request: John Smith (Employee ID: 45832) cannot access his HR portal. Email: john.smith@company.com Issue: Unable to update dependent details for his son born on 12/03/2020. Request: Please resolve urgently." Example of an OOD non-HR ticket (out of scope): "Request: The office coffee machine on the 3rd floor is broken again. Issue: Display shows ’Water Pump Error.’ Request: Please arrange maintenance." Output Format: Return a single JSON object with the following keys: "classification": Either "ID" or "OOD". "reasoning": A one-sentence explanation referencing the rule followed or violated. Task: Analyze the following DOCUMENT and provide your classification: Your JSON response:
@@
LLM Prompt: Policy Compliance Classification (Content-Focused) You are a security/policy analyst. Analyze the user’s query and decide if the content of the INPUT is allowed under the policy. Ignore tooling/functions; focus only on what data the INPUT contains and what the OUTPUT must contain. Policy — Allowed vs. Disallowed Content ALLOWED INPUT CONTENT (no personal data, no secrets) Alert metadata: alert/case IDs, timestamps, severity, source system, rule/signature names, tags. Technical indicators (IoCs): file hashes, public IPs, domains, URLs (no credentials), file/process names, registry keys, event codes. System/asset context: asset IDs, generic hostnames (not derived from real names), service/application names, environment (prod/stage/dev), non-personal ticket/channel IDs. Log details (sanitized): network flow summaries, HTTP metadata (without tokens/cookies), counts, time windows, error codes; fields may be pseudonymized or hashed. Org context (non-personal): role or team names (e.g., “SRE on-call”), runbook references, anonymized user placeholders (user_A, hashed IDs). Threat intel: malware family names, campaign/actor names, rule IDs, confidence scores. Attachments/screenshots only if redacted/pseudonymized so no personal data or secrets are visible. DISALLOWED INPUT CONTENT Personal data (PII): names, emails, phone numbers, home addresses, government IDs, birth dates, employee/customer IDs tied to individuals. Sensitive records: health, financial, or HR records; salary/beneficiary details; customer lists; resumes. Secrets/credentials: passwords, API keys, private keys, tokens, session IDs, cookies, JWTs, MFA codes, OAuth headers. Raw dumps containing PII/secrets: browser/localStorage, memory dumps, unredacted chat/email transcripts, authorization headers. Identifiers revealing a person: hostnames/usernames or filenames that embed real names (must be pseudonymized before submission). Out-of-scope facility issues (e.g., coffee machine maintenance). Output Contract Return a single JSON object: "classification": "ID" if the INPUT only contains Allowed content; otherwise "OOD". "reasoning": One concise sentence citing the specific rule followed/violated. Task: Analyze the following DOCUMENT and classify it: Your JSON response:
@@ A.3 Evaluation for POLICY-TREE-GEN
SOC O1 26.9 22.6 85.7 0.987 100 0.958
SOC GPT-OSS 120B 66.7 10.3 0.0 0.974 0.0 0.881
SOC Llama 405B 40.0 9.8 0.0 0.842 100.0 0.790
SOC Claude 3.5 100.0 5.4 0.0 0.818 100.0 0.782